# Dual Numbers for Arbitrary Order Automatic Differentiation

### v1

@@ Abstract
Dual numbers are a well-known tool for computing derivatives of functions. While the theoretical framework for calculating derivatives of arbitrary order is well established, practical implementations remain less developed. One notable implementation is available in the Julia programming language where dual numbers are designed to be nested, enabling the computation of derivatives to arbitrary order. However, this approach has a significant drawback as it struggles with scalability for high-order derivatives. The nested structure quickly consumes memory, making it challenging to compute derivatives of higher orders. In this study, we introduce DNAOAD, a Fortran-based implementation of automatic differentiation capable of handling derivatives of arbitrary order using dual numbers. This implementation employs a direct approach to represent dual numbers without relying on recursive or nested structures. As a result, DNAOAD facilitates the efficient computation of derivatives of very high orders while addressing the memory limitations of existing methods.
@@ 1 Introduction
Differentiation is a fundamental process that is widely applied in numerous fields of science and engineering. It can be carried out symbolically using specialized software, such as Maxima, Maple, or Mathematica, or numerically through finite-difference methods. Another approach is Automatic Differentiation (AD) [1, 2, 3], which enables the numerical computation of derivatives with the same level of accuracy and precision as symbolic methods, but without requiring symbolic computations.
Automatic Differentiation (AD) is a systematic algorithmic technique for efficiently applying the chain rule to compute derivatives. Although there is extensive research on this topic, most studies focus on the computation of first- and second-order derivatives for real-valued functions [4, 5, 6, 7, 8, 9, 10]. ADOL-C [11, 12, 13] is an interesting implementation of this technique, which takes advantage of operator overloading to enable differentiation. Although ADOL-C supports higher-order derivatives for some functions, a considerable number, including asin, acos, atan, and exponentiation, lack full support for differentiation across all derivative orders.
Notable AD implementations that support the computation of derivatives of arbitrary order include the PyTorch library in Python [14] and the ForwardDiff package in Julia [15]. Even when these libraries do not natively support automatic differentiation for complex-valued functions, the complex case can be managed by separately computing the derivatives of the real and imaginary components of the functions involved.
PyTorch is a versatile and widely used Python library for AD and deep learning applications. One of its key strengths is its ability to compute higher-order derivatives using a dynamic computation graph. This graph tracks operations at runtime and facilitates automatic computation of gradients. However, when dealing with higher-order differentiation—especially when calculating derivatives of complex operations or multiple gradients—PyTorch retains the entire computational graph in memory. While this is necessary to compute successive gradients, it significantly increases memory consumption. For instance, calculating higher-order derivatives can cause the computational graph to rapidly expand due to the accumulation of intermediate values. Consequently, PyTorch’s memory usage may quickly exceed the system’s available memory, often resulting in a std::bad_alloc error, which occurs when memory allocation requests surpass the system’s capacity. In our numerical experiment, performed on a 12th generation Intel(R) Core(TM) i7-12650H CPU @ 4.70 GHz (base 400 MHz) with 15 GB of RAM, running Debian GNU/Linux 12, this error occurred when calculating the 12th-order derivative of the function $f(x)=\sin(x)\,\exp(-x^{2})$, composed with itself five times.
In contrast, ForwardDiff, implemented in Julia, computes derivatives using dual numbers and elegantly handles higher-order derivatives through their recursive nesting. For simplicity, we refer to a structure resulting from nesting these numbers $n$ times as a dual number of order $n$ or a multidual number of order $n$ [16, 17, 18, 19]. Despite the elegance of this approach, challenges emerge when computing higher-order derivatives with ForwardDiff. With each successive derivative, ForwardDiff generates a new Dual type that encapsulates all previously computed derivative information. This recursive nesting increases complexity, as each successive derivative introduces additional layers of dual numbers. For higher-order derivatives, this nesting becomes particularly problematic, leading to excessive memory consumption and ultimately causing stack overflows. In our numerical experiments, a StackOverflowError occurred when attempting to compute the 15th-order derivative of the previously mentioned function.
In this study, we present a Fortran implementation of dual numbers to arbitrary order without relying on a recursive or nested structure. This approach theoretically enables the computation of derivatives of any order, effectively overcoming the limitations of the nested dual number method. While practical constraints, such as computational resources, will ultimately limit the achievable order of derivatives, our method significantly extends the range of computable derivatives beyond what is feasible with the nested approach. It is important to note, however, that as the derivative order increases, the risk of precision loss grows. Despite depending on the specific function being differentiated, our numerical experiments show that derivatives of orders around 100 may require quadruple precision for accurate computation. For even higher-order derivatives, further increases in precision will likely be necessary.
Our Fortran implementation uses complex numbers as the components of dual numbers. Although an implementation with real numbers may be necessary for efficiency reasons, we chose to use complex numbers for greater generality. It is important to note the lack of a consistent convention regarding the branches chosen for complex functions across different programming environments. For example, when computing $\arcsin(1.1)$, the result obtained from SBCL (version 2.2.9.debian), Octave (version 7.3.0), and Wolfram Mathematica (version 12.0) is $(1.5708,-0.44357)$. In contrast, the same computation using Python’s cmath library yields $(1.5708,0.44357)$, a result consistent with that produced by modern versions of Fortran. We have elected to adopt the latter convention, aligning with the Fortran implementation, to avoid the necessity of developing custom functions to select a specific branch. The decision regarding which branch to standardize will ultimately rest with the Fortran standards committee (ISO/IEC JTC1/SC22/WG5).
@@ 2 Dual numbers and derivatives
2.1 First order case
Analogous to the definition of a complex number $z=a+i\,b$ where $a$, $b$ are real numbers and $i$, the imaginary unit satisfying $i^{2}=-1$, a dual numbers is defined as
$r$ $=a_{0}\,\epsilon_{0}+\epsilon_{1}\,a_{1},$ (1)
$r$ $=a_{0}+\epsilon_{1}\,a_{1},$ (2)
where $a_{0}$, $a_{1}$ are real or complex numbers, $\epsilon_{0}=1$, and $\epsilon_{1}$ is the dual unit satisfying
@@ 2.1 First order case
In a similar manner to how an analytical real function is extended to a complex function, we can evaluate the Taylor expansion of a function $f(z)$ at $z+\epsilon_{1}$:
@@
However, in virtue of Eq. (4) $\epsilon_{1}^{k}=0$ for $k>1$. This simplifies the expansion to
@@
Therefore, evaluating an analytic function at the dual number $z+\epsilon_{1}$ yields a dual number where its $\epsilon_{0}$ component is $f(z)$ and its $\epsilon_{1}$ component is $f^{\prime}(z)$. The process of extending $f(z)$ to operate on dual numbers is called dualizing $f(z)$. As a simple example, the dual extension of the sine function $sin(z)$ will be
$sin(z)=\sin z+\cos z\;\epsilon_{1}.$ (7)
Here, the function name in italics indicates the dual version of the original function. It is important to note that this function is not the general dual version for the “$\sin$” function, as $z$ is a complex number. To construct the general dual version we need to be able of computing $sin(g)$ for $g=g_{0}+\epsilon_{1}\,g_{1}$ a dual number. Setting $g(z+\epsilon_{1})$ = $g_{0}+g_{1}\epsilon_{1}$ and evaluating in the Taylor series expansion for the function $f$, we have
$f(g_{0}+\epsilon_{1}\,g_{1})=f(g_{0})+f^{\prime}(g_{0})g_{1}\,\epsilon_{1}.$ (8)
Therefore, the dual extension of the sine function becomes:
$sin(g)=\sin(g_{0})+\cos(g_{0})\,g_{1}\,\epsilon_{1}.$ (9)
This procedure can be applied to dualize all elementary functions [^1] A non-elementary function, such as the error function $\text{erf}(x)$, can be extended to dual numbers in the same way. However, since this function is not implemented in Fortran for complex values, it is not extended in our dual number implementation. and main operators of a programming language.
The generalization for calculating arbitrary-order derivatives is straightforward from a theoretical standpoint. However, computational implementation presents significant challenges, particularly in avoiding stack overflow errors that can occur at higher derivative orders when using recursive or nested structures. Most works in the literature focus on implementing dual numbers for the real case, typically up to the first or second order [20, 21, 22, 23, 19]. Some extensions to third and fourth orders, including the complex case, have also been proposed [16, 17]. While implementations using recursive or nested structures are effective and efficient for lower-order derivatives, they may encounter limitations at very high derivative orders due to stack depth limitations. Achieving dual number implementations for arbitrary orders without relying on recursion is the primary contribution of our study.
2.2 arbitrary order case
Derivatives to arbitrary order can be calculated by defining the dual number [17]
@@
with $a_{k}$ a complex number and
$\epsilon_{i}\cdot\epsilon_{j}=\begin{cases}0&\mbox{if }i+j>n,\vspace*{0.3cm}\\ \frac{(i+j)!}{i!\,j!}\epsilon_{i+j}&\mbox{otherwise. }\end{cases}$ (11)
Now, evaluating the Taylor expansion for an analytical function in the number $z+\epsilon_{1}$ and using Eq.(11) we have
$f(z+\epsilon_{1})=f(z)\epsilon_{0}+f^{\prime}(z)\epsilon_{1}+\dots+f^{(n)}(z)\epsilon_{n}.$ (12)
Computing Eq. (12) for the elementary function is not very complicated, and thus the dual extension for $f(z)$ is relatively easy. However, in the general case, the dual version for $f(g(z))$, i.e. $f(g)$ is much more complicated. Nevertheless, the process is simplified if one first implements the well known Faà Di Bruno formula [24, 25] for dual numbers as it is showed in the next section.
@@ 3 Dual number implementation to arbitrary order
To implement dual numbers to arbitrary order, we will first implement the chain rule to calculate the $n$-th derivative of a function in Fortran. The data structure we will use is shown in Listing 1.
1 type, public :: dualzn
2 complex(prec), allocatable, dimension(:) :: f
3 end type dualzn
Listing 1: Fortran Code for dualzn Type using the precision given by prec. The $k$-th component of a dual number $g$ of this type is accessed by $g\%f(k)$.
Now, according to the famous Faà Di Bruno formula, the $n$-th derivative of $f(g(x))$ is given by
@@
where $D^{n}=d^{n}/dx^{n}$, $f^{(k)}(x)=D^{k}f(x)$ and the the sum is taken over all sequences of nonnegative integers $k_{1},k_{2},\ldots,k_{n}$ such that they obey the Diophantine equation:
@@
Although a recursive formula that avoid solving the Diophantine equation can be obtained [26], a more direct approach is to write Eq. (13) as:
$D^{n}f(g(x))=\sum_{k=1}^{n}f^{(k)}(g(x))\;B_{n,k}(g^{\prime}(x),g^{\prime\prime}(x),\dots,g^{(n-k+1)}(x)),$ (15)
where the $B_{n,k}(x_{1},x_{2},\dots,x_{n-k+1})$ are the partial Bell polynomials which can be recursively defined by [27]
$B_{n,k}(x_{1},x_{2},\dots,x_{n-k+1})=\sum_{i=0}^{n-k}\binom{n-1}{i}x_{i+1}B_{n-i-1,k-1}(x_{1},x_{2},\dots,x_{n-k-i+1}),$ (16)
with
$B_{0,0}(\cdot)$ $=1$ (17)
$B_{n,0}(\cdot)$ $=0\text{~{}for~{}}n\geqslant 1$ (18)
$B_{0,k}(\cdot)$ $=0\text{~{}for~{}}k\geqslant 1.$ (19)
From this recursive definition, we can build the iterative version shown in Algorithm 3.
@@
1:$n,k\in\mathbb{Z}$, $x\in\mathbb{C}^{m}$
2:$result\_value\in\mathbb{C}$
3:Initialize $dp[n+1,k+1]\leftarrow 0$
4:$dp[1,1]\leftarrow 1$
@@
6:     $dp[nn+1,1]\leftarrow 0$
7:end for
8:for $kk\leftarrow 1$ to $k$ do
9:     $dp[1,kk+1]\leftarrow 0$
@@
11:if $n=0$ and $k=0$ then
12:     $result\_value\leftarrow 1$
13:     return $result\_value$
14:else if $\text{size}(x)=0$ then
15:     $result\_value\leftarrow 0$
16:     return $result\_value$
17:end if
18:for $nn\leftarrow 1$ to $n$ do
19:     for $kk\leftarrow 1$ to $k$ do
20:         $LX\leftarrow\text{size}(x)$
21:         $nx\leftarrow\max(nn-kk+1,LX)$
22:         if $nx>0$ then
23:              Allocate $newx[nx]\leftarrow 0$
24:              $newx[1:LX]\leftarrow x$
25:              $sum\_val\leftarrow 0$
26:              for $ii\leftarrow 0$ to $nn-kk$ do
27:                  $sum\_val\leftarrow sum\_val+\binom{nn-1}{ii}\cdot newx[ii+1]\cdot dp[nn-ii,kk]$
28:              end for
29:              $dp[nn+1,kk+1]\leftarrow sum\_val$
30:              Deallocate $newx$
31:         end if
32:     end for
33:end for
34:$result\_value\leftarrow dp[n+1,k+1]$
35:return $result\_value$
Having implemented the $B_{n,k}$ function, we can now implement the chain rule Eq. (15) in Fortran using the derived data type dualzn from Listing 1 and following Algorithm 3. Here, the $fci$ function provides the dual version of $f(z)$. We emphasize that it is not the dual version of $f(g(z))$, which is constructed after applying the chain rule to $f$ through function Dnd. Let us exemplify this with the sinusoidal function.
@@
1:$fci$: Function, $gdual$: dualzn type, $n\in\mathbb{Z}$
2:$dnfcg\in\mathbb{C}$
3:$g0\leftarrow gdual\%f(0)$
4:$fvd\leftarrow fci(g0)$
@@
6:     $dnfcg\leftarrow fvd\%f(0)$
7:else
8:     $suma\leftarrow 0$
@@
10:         Allocate $xvg[1:n-k+1]$
11:         for $j\leftarrow 1$ to $n-k+1$ do
12:              $xvg[j]\leftarrow gdual\%f(j)$
13:         end for
14:         $suma\leftarrow suma+fvd\%f(k)\cdot B(n,k,xvg)$
15:         Deallocate $xvg$
16:     end for
17:     $dnfcg\leftarrow suma$
18:end if
19:return $dnfcg$
The dual version $sin(z)$ to order $n$, of the $\sin(z)$ function is given by
$sin(z)$ $=\sum_{k=0}^{n}D^{k}(\sin z)\,\epsilon_{k}$
$sin(z)$ $=\sum_{k=0}^{n}\sin\left(z+k\,\frac{\pi}{2}\right)\epsilon_{k}.$ (20)
We recall that in this function $sin\;z$ is a dual number but $z$ is not. To construct the dual version $sin(g)$, being $g$ a dual number, we need to apply the Dnd function to $sin\;z$. Therefore
$sin(g)=\sum_{k=0}^{n}\text{Dnd}(sin,g,k)\,\epsilon_{k}$ (21)
will be the dual version of $\sin(g(z))$.
By following this procedure, the dual version for all elementary functions can be obtained. As exemplified in Eq. (20), to obtain the dual version for $f(z)$ its derivatives are necessary. For most of the elementary functions this can be done with no much effort. There are another functions, however, for which to obtain its dual version $f(z)$ could be a little more complicated. Take for instance the inverse of the sine function, the $\arcsin(z)$ function. A recursive [^2] While a recursive approach to calculating the derivatives of functions may appear more direct, we advise against it due to potential inefficiencies, including lower performance and higher memory consumption. formula to calculate its $n$-th derivative is given by
$D^{0}(\arcsin z)$ $=\arcsin z$ (22)
$D^{1}(\arcsin z)$ $=\frac{1}{\sqrt{1-z^{2}}}$ (23)
$D^{n}(\arcsin z)$ $=-\sum_{k=1}^{n-1}\binom{n-1}{k}D^{k}\left(\sqrt{1-z^{2}}\right)\frac{D^{n-k}(\arcsin z)}{\sqrt{1-z^{2}}}.$ (24)
One can pursue this approach; however, if we first dualize the functions $inv(z)=1/z$ as well as the square root function and the difference and multiplication operators, then by expressing Eq. (23) in its dual form, all higher-order derivatives can be automatically computed. All these function are relatively easy to dualize, for instance, Algorithms 3 and 4 shows the dualization of the multiplication operator using the Leibniz rule.
Algorithm 3 Pseudocode for the function timesdzn(A,B,k), the Leibniz rule.
1:$A$: dualzn type, $B$: dualzn type, $k\in\mathbb{Z}$
2:$fr\in\mathbb{C}$
3:$fr\leftarrow 0$

### v2

@@ Abstract
Dual numbers are a well-established tool for computing derivatives and constitute the basis of forward-mode automatic differentiation. While the theoretical framework for computing derivatives of arbitrary order is well understood, practical and scalable implementations remain limited. Existing approaches based on nested dual numbers, such as those used in modern high-level languages, suffer from severe memory growth and poor scalability as the derivative order increases. In this work, we introduce DNAOAD, a Fortran-based automatic differentiation framework capable of computing derivatives of arbitrary order using dual numbers with a direct, non-nested representation. By avoiding recursive data structures, DNAOAD significantly reduces memory usage and enables the efficient computation of derivatives of very high order, overcoming key scalability limitations of existing methods and making it particularly well suited for high-performance scientific computing applications.
@@ 1 Introduction
Differentiation is a fundamental operation in science and engineering, with applications ranging from optimization and sensitivity analysis to numerical methods and scientific modeling. Derivatives can be computed symbolically using computer algebra systems such as Maxima, Maple, or Mathematica, or numerically through finite-difference schemes. While symbolic differentiation provides exact expressions and finite differences are straightforward to implement, both approaches have well-known limitations related to expression growth, numerical stability, and truncation errors. An alternative approach is Automatic Differentiation (AD) [9, 4, 2], which computes derivatives with machine precision by systematically applying the chain rule to numerical programs, without resorting to symbolic manipulation.
Automatic Differentiation is an algorithmic technique that enables the efficient computation of derivatives of functions defined by computer programs. Although AD has been extensively studied, most practical implementations focus on first- and second-order derivatives of real-valued functions [24, 23, 11, 12, 30, 8, 26]. One notable implementation is ADOL-C [10, 29, 1], which employs operator overloading to support differentiation in C/C++. While ADOL-C provides partial support for higher-order derivatives, several elementary functions, including inverse trigonometric functions and exponentiation, are not fully supported across all derivative orders.
Several modern AD frameworks provide mechanisms for computing higher-order derivatives of arbitrary order. Graph-based approaches, such as those used in the PyTorch library [17], rely on dynamic computational graphs that record intermediate operations at runtime. Although this strategy enables repeated differentiation, it requires retaining the entire computational graph in memory, leading to rapid memory growth as the derivative order increases. In practice, this often results in memory exhaustion when computing higher-order derivatives of moderately complex functions.
An alternative strategy is based on forward-mode AD using dual numbers. The ForwardDiff package in Julia [25] supports higher-order derivatives by recursively nesting dual numbers. For simplicity, we refer to a structure obtained by nesting dual numbers $n$ times as a dual number of order $n$, or a multidual number of order $n$ [22, 21, 15, 28]. While this approach is mathematically elegant, the recursive nesting leads to rapid growth in memory usage and computational complexity. As the derivative order increases, the nested structure becomes increasingly inefficient and may ultimately result in stack overflows or memory exhaustion.
In this work, we present a Fortran-based implementation of dual numbers that supports the computation of derivatives of arbitrary order without relying on recursive or nested data structures. By employing a direct representation of dual numbers, the proposed approach avoids the memory explosion inherent to nested methods and significantly extends the range of derivative orders that can be computed in practice. While practical limits are ultimately imposed by available computational resources and numerical precision, the proposed method overcomes key scalability limitations of existing approaches and is particularly well suited for high-performance scientific computing applications. To the authors’ knowledge, this work presents the first practical implementation of dual numbers of arbitrary order that avoids recursive or nested representations while remaining scalable for very high derivative orders.
@@ 2 Dual numbers and derivatives
2.1 First-order case
Analogous to the definition of a complex number z=a+ibz=a+i\,b, where $a,b\in\mathbb{R}$ and $i$ is the imaginary unit satisfying $i^{2}=-1$, a dual number is defined as
$r$ $=a_{0}\,\epsilon_{0}+a_{1}\,\epsilon_{1},$ (1)
$=a_{0}+a_{1}\,\epsilon_{1},$ (2)
where $a_{0}$ and $a_{1}$ are real or complex numbers, $\epsilon_{0}=1$, and $\epsilon_{1}$ is the dual unit satisfying
@@ 2.1 First-order case
In a manner analogous to extending a real function to the complex domain, an analytic function $f(z)$ can be evaluated at a dual argument $z+\epsilon_{1}$ by means of its Taylor expansion,
@@
However, due to Eq. (3), all powers $\epsilon_{1}^{k}$ vanish for $k>1$, and the expansion reduces to
@@
Thus, evaluating an analytic function at the dual number $z+\epsilon_{1}$ produces a dual number whose $\epsilon_{0}$ component is $f(z)$ and whose $\epsilon_{1}$ component is $f^{\prime}(z)$. The extension of $f(z)$ to operate on dual numbers is commonly referred to as dualizing the function.
As a simple example, the dual extension of the sine function evaluated at $z+\epsilon_{1}$ is
$\sin(z+\epsilon_{1})=\sin z+\cos z\,\epsilon_{1}.$ (6)
Here, the italicized function name denotes the dual version of the original function. This expression, however, corresponds only to the special case in which the argument is $z+\epsilon_{1}$, with $z$ a complex number. To construct the general dual extension, consider a dual number $g=g_{0}+g_{1}\,\epsilon_{1}$. Substituting this expression into the Taylor expansion of an analytic function $f$, we obtain
$f(g_{0}+g_{1}\,\epsilon_{1})=f(g_{0})+f^{\prime}(g_{0})\,g_{1}\,\epsilon_{1}.$ (7)
Accordingly, the general dual extension of the sine function is given by
$\sin(g)=\sin(g_{0})+\cos(g_{0})\,g_{1}\,\epsilon_{1}.$ (8)
This procedure can be applied to dualize all elementary functions and the main operators of a programming language [^1] Non-elementary functions, such as the error function $\mathrm{erf}(x)$, can in principle be extended to dual numbers in the same manner. However, since this function is not implemented for complex arguments in Fortran, it is not included in the present implementation..
From a theoretical perspective, the generalization of this approach to higher-order derivatives is straightforward. In practice, however, computational implementation becomes increasingly challenging, particularly when recursive or nested data structures are employed. Most existing implementations focus on real-valued dual numbers and are restricted to first- or second-order derivatives [5, 7, 31, 19, 28]. Extensions to third- and fourth-order derivatives, including the complex case, have also been proposed [22, 21]. While recursive and nested representations are effective for low derivative orders, they may encounter severe limitations at higher orders due to excessive memory usage and stack depth constraints. Eliminating these limitations by avoiding recursion is a central motivation of the present work.
2.2 Arbitrary-order case
Derivatives of arbitrary order can be computed by defining a dual number of order $n$ as [21]
@@ 2.2 Arbitrary-order case
where $a_{k}$ are complex coefficients and the basis elements $\epsilon_{k}$ satisfy the multiplication rule
$\epsilon_{i}\cdot\epsilon_{j}=\begin{cases}0,&\text{if }i+j>n,\\[5.69046pt] \dfrac{(i+j)!}{i!\,j!}\,\epsilon_{i+j},&\text{otherwise}.\end{cases}$ (10)
Evaluating the Taylor expansion of an analytic function at $z+\epsilon_{1}$ and using Eq. (10) yields
$f(z+\epsilon_{1})=f(z)\,\epsilon_{0}+f^{\prime}(z)\,\epsilon_{1}+\cdots+f^{(n)}(z)\,\epsilon_{n}.$ (11)
The dual extension of elementary functions can therefore be constructed directly. In the general case of a composite function $f(g(z))$, the corresponding dual extension $f(g)$ becomes significantly more involved. This difficulty can be addressed by implementing the Faà di Bruno formula [13, 6] within the dual number framework, as discussed in the following section.
@@ 3 Dual number implementation to arbitrary order
To implement dual numbers of arbitrary order, we first construct a non-recursive formulation of the chain rule suitable for numerical evaluation in Fortran. The core data structure employed throughout this work is the derived type shown in Listing 1, which stores the coefficients of a dual number of order $n$ in a one-dimensional array.
1type, public :: dualzn
2 complex(prec), allocatable :: f(:)
3end type dualzn
Listing 1: Fortran derived type dualzn using the precision specified by prec. The $k$-th coefficient of a dual number $g$ is accessed as g%f(k).
This representation avoids recursive or nested data structures and allows all derivative components to be accessed directly by index, which is essential for scalability at high derivative orders.
3.1 Chain rule via Faà di Bruno formula
According to the Faà di Bruno formula, the $n$-th derivative of a composite function $f(g(x))$ is given by
@@ 3.1 Chain rule via Faà di Bruno formula
where $D^{n}=d^{n}/dx^{n}$, $f^{(k)}(x)=D^{k}f(x)$, and the sum is taken over all nonnegative integer solutions of the Diophantine equation
@@
Although recursive formulations that avoid explicitly solving this Diophantine equation exist [16], a more convenient and computationally efficient formulation is obtained by rewriting Eq. (12) as
$D^{n}f(g(x))=\sum_{k=1}^{n}f^{(k)}(g(x))\,B_{n,k}\!\left(g^{\prime}(x),g^{\prime\prime}(x),\dots,g^{(n-k+1)}(x)\right),$ (14)
where $B_{n,k}$ are the partial Bell polynomials.
The partial Bell polynomials can be defined recursively as [3]
$B_{n,k}(x_{1},\dots,x_{n-k+1})$ $=\sum_{i=0}^{n-k}\binom{n-1}{i}\,x_{i+1}\,B_{n-i-1,k-1}(x_{1},\dots,x_{n-k-i+1}),$ (15)
with initial conditions
$B_{0,0}$ $=1,$ (16)
$B_{n,0}$ $=0\quad(n\geq 1),$ (17)
$B_{0,k}$ $=0\quad(k\geq 1).$ (18)
From this definition, an iterative dynamic-programming implementation can be constructed. Algorithm 3.1 presents pseudocode for the evaluation of $B_{n,k}$.
@@
1 Iterative computation of the partial Bell polynomial $B_{n,k}(x)$.
[…]
3:Initialize $dp[0:n,0:k]\leftarrow 0$
[…]
@@
6:  for $kk\leftarrow 1$ to $\min(nn,k)$ do
7:   for $i\leftarrow 0$ to nnkknn-$k$ do
8:     $dp[nn,kk]\leftarrow dp[nn,kk]+\binom{nn-1}{i}\,x_{i+1}\,dp[nn-i-1,kk-1]$
9:   end for
@@
11:end for
12:return $dp[n,k]$
3.2 Implementation of the chain rule
Once the partial Bell polynomials are available, the chain rule (14) can be implemented directly for dual numbers of type dualzn. Algorithm 3.2 shows the pseudocode for the function Dnd, which computes the $n$-th derivative of $f(g(x))$. The function fci provides the dual extension of $f(z)$ evaluated at a scalar argument.
@@ 3.2 Implementation of the chain rule
2 Pseudocode for Dnd(fci, gdual, n) implementing the chain rule.
1:fci: function, gdual: dualzn, $n\in\mathbb{Z}$
[…]
@@
6:  return $fvd\%f(0)$
7:end if
[…]
@@
10:  $sum\leftarrow sum+fvd\%f(k)\,B_{n,k}(gdual\%f(1:n-k+1))$
11:end for
12:return sumsum
3.3 Dualization of elementary functions
As an illustrative example, the dual extension of $\sin(z)$ to order $n$ is
$\sin(z)$ $=\sum_{k=0}^{n}D^{k}(\sin z)\,\epsilon_{k}=\sum_{k=0}^{n}\sin\!\left(z+k\frac{\pi}{2}\right)\epsilon_{k}.$ (19)
Here, $\sin(z)$ denotes a dual number whose argument $z$ is scalar. To construct the dual extension $\sin(g)$, where $g$ is a dual number, the chain rule must be applied:
$\sin(g)=\sum_{k=0}^{n}\texttt{Dnd}(\sin,g,k)\,\epsilon_{k}.$ (20)
Following this procedure, dual extensions for all elementary functions can be constructed. In cases where closed-form expressions for higher derivatives are cumbersome, functions can be dualized by combining previously dualized elementary operations. For example, although the higher derivatives of $\arcsin(z)$ satisfy the recursive relations
$D^{0}(\arcsin z)$ $=\arcsin z,$ (21)
$D^{1}(\arcsin z)$ $=\frac{1}{\sqrt{1-z^{2}}},$ (22)
$D^{n}(\arcsin z)$ $=-\sum_{k=1}^{n-1}\binom{n-1}{k}D^{k}\!\left(\sqrt{1-z^{2}}\right)\frac{D^{n-k}(\arcsin z)}{\sqrt{1-z^{2}}},$ (23)
it is computationally preferable to dualize the inverse, square root, multiplication, and subtraction operators, and to rely on the automatic propagation of derivatives through these operations.
As an example, Algorithms 3 and 4 show the dualization of multiplication using the Leibniz rule.
Algorithm 3 Leibniz rule for the $k$-th derivative of a product.
1:$A,B$: dualzn, $k\in\mathbb{Z}$
[…]